### v1

@@ 1 Introduction
The skin, the largest organ of the human body, protects the internal organs from external hazards such as dust, heat, polluted water, and ultraviolet rays. However, the primary line of defense is frequently exposed to external agents causing several skin diseases including cancer [1]. Over the past few years, the number of skin cancer patients has grown significantly, accounting for one-third of the reported cancers [2, 3]. Benign and malignant skin cancers are two distinct types with contrasting characteristics [4]. Benign skin cancers typically have a modest growth rate and remain confined, without spreading to other tissues. They are usually not life-threatening, though they could be uncomfortable if left untreated [5]. Malignant skin cancers, on the other hand, are more aggressive in their nature. They could infiltrate adjacent tissues and, if they are not found and treated promptly, spread to other areas of the body, so becoming a major health hazard [6]. Skin tumors that are malignant frequently have uneven borders and proliferate quickly. Leveraging computer vision-based technologies for the early detection of skin cancer offers the potential for enhanced efficiency and accuracy compared to traditional approaches.
@@
Traditional image classification techniques primarily rely on Convolutional Neural Networks (CNNs). Convolutional layers and the pooling layers are the essential elements of CNNs. While convolutional layers effectively extract useful features from the input image, the pooling layer downsamples the spatial dimensions of the feature map, conserving the important features and enhancing computational efficiency [7]. Although the CNNs excel in a variety of image recognition tasks, the networks require high computational power and a large number of labeled training data which is often expensive to obtain. Transfer learning offers a solution to this challenge by enabling the utilization of pretrained neural networks trained on extensive datasets containing generic images. Throughout the training process, the model learns to extract useful features that can be leveraged in the related tasks. Since a new task begins with a pretrained CNN backbone, it requires only the top layers to be fine-tuned which significantly reduces the computational burden [8]. Furthermore, transfer learning also allows the model to work with less training data since only a small section of the model needs to be trained.
@@
Another restriction of CNNs is their limited ability to accommodate global context. Since CNNs incorporate convolutional layers that are limited to their kernel size, the networks fail to capture larger and more global patterns beyond the receptive field of the kernel. Vision Transformer (ViT) [9] has emerged as a solution to this problem. The ViT, originated from the transformer [10] architecture, incorporates self-attention mechanism. Dissimilar to CNNs, ViT’s self-attention mechanism captures long-range dependencies and global context from the input image. The transformer devices an input image into non-overlapping segments named patches which are treated as sequence tokens and processed using self-attention layers. By doing so, ViT overcomes CNNs’ shortcomings in handling global context, making it especially useful for applications that require comprehending relationships across all segments are essential. The use of Vision Transformer has yielded encouraging results in a variety of computer vision applications, outperforming state-of-the-art CNN architectures [11, 12, 13].
@@
The growing size and complexity of deep learning models present a substantial challenge, especially when deployed on resource-constrained devices. As these models grow in size, their computational and memory needs increase, making them unsuitable for practical usage on devices with limited capabilities, such as mobile phones or edge devices [14, 15]. To solve this issue, a mechanism has arisen for compressing these bulky models into more efficient alternatives while maintaining performance. This procedure entails training a smaller model, commonly known as the student model, to match the predictions of a bigger, pre-existing model, known as the teacher model. The student model learns from both labeled data and the knowledge embedded in the softer probability distributions provided by the teacher model [16]. This information transmission not only allows for the design of more compact models but also improves their generalizability [17]. The ability to deploy these streamlined models on resource-constrained devices makes deep learning applications more accessible and useful in real-world scenarios.
@@ 2 Related Works
Lately, considerable attention has been directed towards achieving high-accuracy classification of skin cancer. Among them, the CNN-based methods have emerged as the most frequently employed techniques. Ali et al. [18] investigated eight EffecientNets (B0-B7) for effectively classifying skin cancer. The proposed approach accommodated an image preprocessing pipeline that consisted of three sequential stages: hair removal, data augmentation, and image resizing. Following the image preprocessing step, they fine-tuned the EffecientNets with different hyper-parameters. While fine-tuning EffecieneNet B0-B5 only required training the top layers, made of fully connected neural networks, the training process of EffecientNet B6-B7 incorporated two steps. In the first step, the top layer was trained while keeping the convolutional feature extractors frozen. In the second step, the last four convolutional blocks were trained, keeping the rest of the model frozen. Moreover, EffecientNet B0-B5 were trained with Stochastic Gradient Descent (SGD) optimizer while EffecientNet B6 and B7 were trained on Adam optimizer since Adam produced better performance on the larger models. Among the eight EfficientNet models, EffecientNet B4 produced the best performance with an accuracy of over 87%. This research work, however, does not present any ablation study leaving the impact of the hyperparameters unknown. Mridha et al. [19] presented a lightweight model having over 5 million parameters. The model adeptly extracted features by employing only two convolutional blocks, each consisting of two convolutional layers along with one max pooling layer and a dropout layer. The system incorporated Grad-CAM and Grad-CAM++ for visualizing the attention heatmaps. Nevertheless, this model failed to achieve a higher accuracy due to its lightweight nature. Zia et al. [2] experimented with MobileNetV2 and DenseNet201, two lightweight image classifiers. In that research, the authors presented two modified classifiers with MobileNetV2 and DenseNet201 backbone which followed three additional convolutional layers for extracting additional features. According to the experiment, the fine-tuned models significantly outperformed the base models. Moreover, the performance of DenseNet201 exceeded MobileNetV2. The authors also employed Grad-CAM for visualizing the attention heatmap of the classifier. This research, however, only leveraged one dataset. Gururaj et al. [20] employed various image preprocessing techniques along with two state-of-the-art image classifiers for classifying skin cancer images from HAM-1000 dataset. The image preprocessing techniques included both oversampling and undersampling to address the data imbalance problem, Dull Razor method for removing noise, and segmentation of the lesion regions with an autoencoder. Finally, the images were classified using DenseNet169 and ResNet50. The best performance was achieved by DenseNet169 with undersampling. Since the system required an image to pass through two CNN models (i.e., autoencoder and DenseNet169), the computational complexity is very high.
@@
Recently transformer-based models have gained attention due to their high classification performance. Xin et al. [21] modified Vision Transformer (ViT) by integrating multiscale an overlapping sliding window. Furthermore, the improved transformer model incorporated multi-scale patch embedding. The proposed ViT was evaluated on HAM1000 and a custom dataset where the model achieved over 94% accuracy. Despite the sliding window mechanism, the self-attention process consumed very high computational resources making the model slow for high-resolution images. Ayas [22] leveraged Swin Transformer for effectively classifying skin cancer images. Swin Transformer [23] is a modified Vision Transformer capable of producing high classification accuracy with the help of shifted window mechanism along with hierarchical feature extraction. According to the research, the Swin Transformer was able to achieve an outstanding performance in differentiating eight types of skin cancers. While there were smaller Swin Transformer models available, the best performance was produced by Large_22K which was responsible for high resource consumption. Since the feature extraction process of CNN and transformer is different, combining these two results in more non-overlapping features which leads to higher performance. Therefore, Hao et al. [24] proposed a fusion model named ConvNeXt-ST-AFF that fuses a CNN model (ConvNeXt) and a transformer model (Swin Transformer). This architecture effectively extracts useful features from two state-of-the-art models and combines them for final prediction. Similar to the previous models, this achieves a high performance by consuming high resources.
@@
Since single model classifiers sometimes fail to achieve significant accuracy, ensembling multiple classifiers produces better performance. Imran et al. [25] fused three deep learning models namely CapsNet, VGG and ResNet for constructing an enable model. Although the model achieved an outstanding performance of 93.5% accuracy, surpassing recent works, its heavyweight nature made it an impractical option for low-constraint devices. Keerthana et al. [26] introduced a fusion model composed of two pre-trained deep neural networks. The features extracted from these networks were concatenated and subsequently fed into a machine learning model for the final classification. Hasan et al. [27] presented a stacked ensemble method named dermoexpert. This model extracted features using three CNN models. In the subsequent stage, the features from the first two CNN streams were fused to form a fourth stream. The fusion process included fusion by channel-averaging and fusion by channel concatenation. Following the second stage, a series of fully connected blocks were integrated with the streams. Finally, the prediction of the four streams was combined using soft voting for the final classification. Similar to the previous works, this research also consumed very high computational resources.
@@
Ali et al. [18] HAM10000 A comparison is made among eight EffecientNets. No ablation study presented.
Mridha et al. [19] HAM10000 Presented an interpretable model. Relatively lower classification accuracy.
Zia et al. [2] Kaggle Presents a novel fine-tuning approach of MobileNet and DenseNet. Leverages only one dataset.
Gururaj et al. [20] HAM10000 Very high classification performance. Computational complexity is also high.
Xin et al. [21] HAM10000, Private dataset A novel transformer architecture with very high performance. The model is very slow for high-resolution images.
Ayas [22] ISIC 2019 Achieves very high performance. Consumes high computation.
Hao et al. [24] ISIC 2019, Private dataset Presents a novel architecture made of state-of-the-art models. Consumes high computation.
Imran et al. [25] ISIC A fusion model with very high performance The heavyweight nature makes it difficult to use for low-constraint devices.
Keerthana et al. [26] ISBI 2016 Presents a detailed comparison of state-of-the-art image classifiers Consumes high computational resources.
Hasan et al. [27] ISIC 2016, ISIC 2017, ISIC 2018 Presented a stacked enseble model that achieved very high classification performance. Very high computaional resource is required.
@@ 3.1 Dataset Details
In the evaluation of the proposed model, we have utilized two benchmark datasets. The first is the Human Against Machine 10000 dataset, commonly referred to as HAM10000 [28], and the second is the Kaggle dataset [29]. The HAM10000 consists of 10,015 images distributed across seven classes. We have categorized these seven classes into two groups: benign and malignant for classification. In the dataset, classes such as Melanoma, Basal cell carcinoma, Actinic keratosis, Benign keratosis-like lesions, and Vascular lesions are categorized as malignant. On the other hand, Melanocytic nevi and Dermatofibroma are classified as benign. Notably, the Kaggle dataset, encompassing 3,297 images, already includes pre-categorized images as benign or malignant. While the images in the HAM10000 dataset are of size 450$\times$600, the images in the Kaggle dataset are already resized to 224$\times$224. Table 2 presents a class-wise distribution of the number of images. For model training and evaluation, we have divided the dataset into a ratio of 70:10:20 for training, validation, and testing, respectively.
@@ 3.2.1 Morphological Closing
Since skin cancer images contain unnecessary hairs which often hinder the classification process, morphological closing is applied at the very first stage of the image preprocessing. Morphological closing is a mathematical operation that involves dilation followed by erosion. It removes small gaps while preserving the original structure in an image [30]. The morphological closing operation can be denoted by equation 1. In this equation, $I$ denotes the input image and $B$ is the structuring element. A structuring element is a kernel that guides neighborhood operations like dilation and erosion. For the proposed method, a structuring element of size 5$\times$5 is selected to strike a balance between the efficiency of hair removal and the risk of leaving behind small remnants. The higher kernel size ensures the removal of moderately sized hair strands. In this operation, the input image is first dilated with the structural element. The dilation process, denoted by $\oplus$, causes objects in the image to expand in size, filling in small gaps. Subsequently, erosion, denoted by $\ominus$, is applied to the resulting image for shrinking the boundaries of regions.
@@ 3.2.2 Bilateral Filtering
Skin cancer images typically contain noise that is challenging to eliminate with conventional low-pass filters, since they may remove subtle diagnostic features. Therefore, we have employed bilateral filter to remove noise while preserving fine details. Unlike standard smoothing filters, which blur the entire image uniformly, bilateral filtering takes into account both the spatial distance between pixels and the intensity difference [31, 32]. It operates by averaging the intensity of surrounding pixels and assigning additional weights to pixels that are spatially close and have similar intensities. The mathematical equation of bilateral filtering is expressed in Equation 2. For an input image $I$, the filtered intensity value at position $q$ can be obtained by considering both the spatial distance, $\|p-q\|$, and the intensity differences, $|I_{p}-I_{q}|$ from the central pixel $p$ and its neighborhood pixels $q$ in the spatial neighborhood $S$. While measuring the distance, the spatial Gaussian filter, $G_{\sigma_{s}}$, and the range Gaussian filter, $G_{\sigma_{r}}$ ensures that neighboring pixels with comparable intensity have a higher influence on the filtered image. Finally, the output is normalized with $W_{p}$ to prevent over-amplification.

### v2

@@ 1 Introduction
The skin, the largest organ of the human body, protects the internal organs from external hazards such as dust, heat, polluted water, and ultraviolet rays. However, the primary line of defense is frequently exposed to external agents causing several skin diseases including cancer [roger2022neuroimmune]. Over the past few years, the number of skin cancer patients has grown significantly, accounting for one-third of the reported cancers [zia2022classification, aladhadh2022effective]. Benign and malignant skin cancers are two distinct types with contrasting characteristics [naqvi2022benign]. Benign skin cancers typically have a modest growth rate and remain confined, without spreading to other tissues. They are usually not life-threatening, though they could be uncomfortable if left untreated [gouda2022detection]. Malignant skin cancers, on the other hand, are more aggressive in their nature. They could infiltrate adjacent tissues and, if they are not found and treated promptly, spread to other areas of the body, so becoming a major health hazard [bratchenko2022classification]. Skin tumors that are malignant frequently have uneven borders and proliferate quickly. Leveraging computer vision-based technologies for the early detection of skin cancer offers the potential for enhanced efficiency and accuracy compared to traditional approaches.
@@
Traditional image classification techniques primarily rely on Convolutional Neural Networks (CNNs). Convolutional layers and the pooling layers are the essential elements of CNNs. While convolutional layers effectively extract useful features from the input image, the pooling layer downsamples the spatial dimensions of the feature map, conserving the important features and enhancing computational efficiency [pi2022broadband]. Although the CNNs excel in a variety of image recognition tasks, the networks require high computational power and a large number of labeled training data which is often expensive to obtain. Transfer learning offers a solution to this challenge by enabling the utilization of pretrained neural networks trained on extensive datasets containing generic images. Throughout the training process, the model learns to extract useful features that can be leveraged in the related tasks. Since a new task begins with a pretrained CNN backbone, it requires only the top layers to be fine-tuned which significantly reduces the computational burden [iman2023review]. Furthermore, transfer learning also allows the model to work with less training data since only a small section of the model needs to be trained.
@@
Another restriction of CNNs is their limited ability to accommodate global context. Since CNNs incorporate convolutional layers that are limited to their kernel size, the networks fail to capture larger and more global patterns beyond the receptive field of the kernel. Vision Transformer (ViT) [dosovitskiy2020image] has emerged as a solution to this problem. The ViT, originated from the transformer [vaswani2017attention] architecture, incorporates self-attention mechanism. Dissimilar to CNNs, ViT’s self-attention mechanism captures long-range dependencies and global context from the input image. The transformer devices an input image into non-overlapping segments named patches which are treated as sequence tokens and processed using self-attention layers. By doing so, ViT overcomes CNNs’ shortcomings in handling global context, making it especially useful for applications that require comprehending relationships across all segments are essential. The use of Vision Transformer has yielded encouraging results in a variety of computer vision applications, outperforming state-of-the-art CNN architectures [shamshad2023transformers, amjoud2023object, thisanke2023semantic].
@@
The growing size and complexity of deep learning models present a substantial challenge, especially when deployed on resource-constrained devices. As these models grow in size, their computational and memory needs increase, making them unsuitable for practical usage on devices with limited capabilities, such as mobile phones or edge devices [islam2023toward, beyer2022knowledge]. To solve this issue, a mechanism has arisen for compressing these bulky models into more efficient alternatives while maintaining performance. This procedure entails training a smaller model, commonly known as the student model, to match the predictions of a bigger, pre-existing model, known as the teacher model. The student model learns from both labeled data and the knowledge embedded in the softer probability distributions provided by the teacher model [song2022spot]. This information transmission not only allows for the design of more compact models but also improves their generalizability [xu2023teacher]. The ability to deploy these streamlined models on resource-constrained devices makes deep learning applications more accessible and useful in real-world scenarios.
@@ 2 Related Works
Lately, considerable attention has been directed towards achieving high-accuracy classification of skin cancer. Among them, the CNN-based methods have emerged as the most frequently employed techniques. Ali et al. [ali2022multiclass] investigated eight EffecientNets (B0-B7) for effectively classifying skin cancer. The proposed approach accommodated an image preprocessing pipeline that consisted of three sequential stages: hair removal, data augmentation, and image resizing. Following the image preprocessing step, they fine-tuned the EffecientNets with different hyper-parameters. While fine-tuning EffecieneNet B0-B5 only required training the top layers, made of fully connected neural networks, the training process of EffecientNet B6-B7 incorporated two steps. In the first step, the top layer was trained while keeping the convolutional feature extractors frozen. In the second step, the last four convolutional blocks were trained, keeping the rest of the model frozen. Moreover, EffecientNet B0-B5 were trained with Stochastic Gradient Descent (SGD) optimizer while EffecientNet B6 and B7 were trained on Adam optimizer since Adam produced better performance on the larger models. Among the eight EfficientNet models, EffecientNet B4 produced the best performance with an accuracy of over 87%. This research work, however, does not present any ablation study leaving the impact of the hyperparameters unknown. Mridha et al. [mridha2023interpretable] presented a lightweight model having over 5 million parameters. The model adeptly extracted features by employing only two convolutional blocks, each consisting of two convolutional layers along with one max pooling layer and a dropout layer. The system incorporated Grad-CAM and Grad-CAM++ for visualizing the attention heatmaps. Nevertheless, this model failed to achieve a higher accuracy due to its lightweight nature. Zia et al. [zia2022classification] experimented with MobileNetV2 and DenseNet201, two lightweight image classifiers. In that research, the authors presented two modified classifiers with MobileNetV2 and DenseNet201 backbone which followed three additional convolutional layers for extracting additional features. According to the experiment, the fine-tuned models significantly outperformed the base models. Moreover, the performance of DenseNet201 exceeded MobileNetV2. The authors also employed Grad-CAM for visualizing the attention heatmap of the classifier. This research, however, only leveraged one dataset. Gururaj et al. [gururaj2023deepskin] employed various image preprocessing techniques along with two state-of-the-art image classifiers for classifying skin cancer images from HAM-1000 dataset. The image preprocessing techniques included both oversampling and undersampling to address the data imbalance problem, Dull Razor method for removing noise, and segmentation of the lesion regions with an autoencoder. Finally, the images were classified using DenseNet169 and ResNet50. The best performance was achieved by DenseNet169 with undersampling. Since the system required an image to pass through two CNN models (i.e., autoencoder and DenseNet169), the computational complexity is very high.
@@
Recently transformer-based models have gained attention due to their high classification performance. Xin et al. [xin2022improved] modified Vision Transformer (ViT) by integrating multiscale an overlapping sliding window. Furthermore, the improved transformer model incorporated multi-scale patch embedding. The proposed ViT was evaluated on HAM1000 and a custom dataset where the model achieved over 94% accuracy. Despite the sliding window mechanism, the self-attention process consumed very high computational resources making the model slow for high-resolution images. Ayas [ayas2023multiclass] leveraged Swin Transformer for effectively classifying skin cancer images. Swin Transformer [liu2021swin] is a modified Vision Transformer capable of producing high classification accuracy with the help of shifted window mechanism along with hierarchical feature extraction. According to the research, the Swin Transformer was able to achieve an outstanding performance in differentiating eight types of skin cancers. While there were smaller Swin Transformer models available, the best performance was produced by Large_22K which was responsible for high resource consumption. Since the feature extraction process of CNN and transformer is different, combining these two results in more non-overlapping features which leads to higher performance. Therefore, Hao et al. [hao2023convnext] proposed a fusion model named ConvNeXt-ST-AFF that fuses a CNN model (ConvNeXt) and a transformer model (Swin Transformer). This architecture effectively extracts useful features from two state-of-the-art models and combines them for final prediction. Similar to the previous models, this achieves a high performance by consuming high resources.
@@
Since single model classifiers sometimes fail to achieve significant accuracy, ensembling multiple classifiers produces better performance. Imran et al. [imran2022skin] fused three deep learning models namely CapsNet, VGG and ResNet for constructing an enable model. Although the model achieved an outstanding performance of 93.5% accuracy, surpassing recent works, its heavyweight nature made it an impractical option for low-constraint devices. Keerthana et al. [keerthana2023hybrid] introduced a fusion model composed of two pre-trained deep neural networks. The features extracted from these networks were concatenated and subsequently fed into a machine learning model for the final classification. Hasan et al. [hasan2022dermoexpert] presented a stacked ensemble method named dermoexpert. This model extracted features using three CNN models. In the subsequent stage, the features from the first two CNN streams were fused to form a fourth stream. The fusion process included fusion by channel-averaging and fusion by channel concatenation. Following the second stage, a series of fully connected blocks were integrated with the streams. Finally, the prediction of the four streams was combined using soft voting for the final classification. Similar to the previous works, this research also consumed very high computational resources.
@@
Ali et al. [ali2022multiclass] HAM10000 A comparison is made among eight EffecientNets. No ablation study presented.
Mridha et al. [mridha2023interpretable] HAM10000 Presented an interpretable model. Relatively lower classification accuracy.
Zia et al. [zia2022classification] Kaggle Presents a novel fine-tuning approach of MobileNet and DenseNet. Leverages only one dataset.
Gururaj et al. [gururaj2023deepskin] HAM10000 Very high classification performance. Computational complexity is also high.
Xin et al. [xin2022improved] HAM10000, Private dataset A novel transformer architecture with very high performance. The model is very slow for high-resolution images.
Ayas [ayas2023multiclass] ISIC 2019 Achieves very high performance. Consumes high computation.
Hao et al. [hao2023convnext] ISIC 2019, Private dataset Presents a novel architecture made of state-of-the-art models. Consumes high computation.
Imran et al. [imran2022skin] ISIC A fusion model with very high performance The heavyweight nature makes it difficult to use for low-constraint devices.
Keerthana et al. [keerthana2023hybrid] ISBI 2016 Presents a detailed comparison of state-of-the-art image classifiers Consumes high computational resources.
Hasan et al. [hasan2022dermoexpert] ISIC 2016, ISIC 2017, ISIC 2018 Presented a stacked enseble model that achieved very high classification performance. Very high computaional resource is required.
@@ 3.1 Dataset Details
In the evaluation of the proposed model, we have utilized two benchmark datasets. The first is the Human Against Machine 10000 dataset, commonly referred to as HAM10000 [tschandl2018ham10000], and the second is the Kaggle dataset [kaggle_skin_cancer]. The HAM10000 consists of 10,015 images distributed across seven classes. We have categorized these seven classes into two groups: benign and malignant for classification. In the dataset, classes such as Melanoma, Basal cell carcinoma, Actinic keratosis, Benign keratosis-like lesions, and Vascular lesions are categorized as malignant. On the other hand, Melanocytic nevi and Dermatofibroma are classified as benign. Notably, the Kaggle dataset, encompassing 3,297 images, already includes pre-categorized images as benign or malignant. While the images in the HAM10000 dataset are of size 450$\times$600, the images in the Kaggle dataset are already resized to 224$\times$224. Table 2 presents a class-wise distribution of the number of images. For model training and evaluation, we have divided the dataset into a ratio of 70:10:20 for training, validation, and testing, respectively.
@@ 3.2.1 Morphological Closing
Since skin cancer images contain unnecessary hairs which often hinder the classification process, morphological closing is applied at the very first stage of the image preprocessing. Morphological closing is a mathematical operation that involves dilation followed by erosion. It removes small gaps while preserving the original structure in an image [lu2021data]. The morphological closing operation can be denoted by equation 1. In this equation, $I$ denotes the input image and $B$ is the structuring element. A structuring element is a kernel that guides neighborhood operations like dilation and erosion. For the proposed method, a structuring element of size 5$\times$5 is selected to strike a balance between the efficiency of hair removal and the risk of leaving behind small remnants. The higher kernel size ensures the removal of moderately sized hair strands. In this operation, the input image is first dilated with the structural element. The dilation process, denoted by $\oplus$, causes objects in the image to expand in size, filling in small gaps. Subsequently, erosion, denoted by $\ominus$, is applied to the resulting image for shrinking the boundaries of regions.
@@ 3.2.2 Bilateral Filtering
Skin cancer images typically contain noise that is challenging to eliminate with conventional low-pass filters, since they may remove subtle diagnostic features. Therefore, we have employed bilateral filter to remove noise while preserving fine details. Unlike standard smoothing filters, which blur the entire image uniformly, bilateral filtering takes into account both the spatial distance between pixels and the intensity difference [li2021multimodal, khattar2022computer]. It operates by averaging the intensity of surrounding pixels and assigning additional weights to pixels that are spatially close and have similar intensities. The mathematical equation of bilateral filtering is expressed in Equation 2. For an input image $I$, the filtered intensity value at position $q$ can be obtained by considering both the spatial distance, $\|p-q\|$, and the intensity differences, $|I_{p}-I_{q}|$ from the central pixel $p$ and its neighborhood pixels $q$ in the spatial neighborhood $S$. While measuring the distance, the spatial Gaussian filter, $G_{\sigma_{s}}$, and the range Gaussian filter, $G_{\sigma_{r}}$ ensures that neighboring pixels with comparable intensity have a higher influence on the filtered image. Finally, the output is normalized with $W_{p}$ to prevent over-amplification.